# Multilevel Markov Chain Monte Carlo with likelihood scaling

### v2

@@ 1 Introduction
(1) $u^{\text{obs}}=\mathcal{F}(\vartheta)+\eta$
where $\eta$ represents the observational noise. The specific structure of $\eta$ impacts the observations and care should be taken when modelling this noise. The noise model used in this paper will be specified in Section 2. A common example of model 1 is when $\mathcal{F}$ is the projection onto the data space of the solution to some known partial differential equation (PDE), and $\vartheta$ are some unknown input parameters to the PDE. Within the context of this paper, we will consider a 2D plane stress problem, where the equations of linear elasticity are solved using linear finite elements (FE). These equations are introduced in Section 2.
@@ 2.1 Forward problem
(3) $\tilde{\lambda}=\dfrac{\vartheta\tilde{\nu}}{(1+\tilde{\nu})(1-2\tilde{\nu})},\qquad\tilde{\mu}=\frac{\vartheta}{2(1+\tilde{\nu})}$
@@
(4) $-\nabla\cdot\left[\tilde{\lambda}(\nabla\cdot u)I+\tilde{\mu}(\nabla u+(\nabla u)^{T})\right]=F_{\text{body}}$
In this equation, the term $F_{\text{body}}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ contains the body forces acting in the system, and $I$ denotes the identity matrix. Equation 4 links the spatially varying Young’s modulus to the displacement field $u:\mathbb{R}^{2}\to\mathbb{R}^{2}$. For practical calculations, we discretize this PDE on a finite element grid. We consider linear rectangular elements for the discretization of the underlying parameters [16].
@@ 2.2 High-resolution observations
(5) $\tilde{u}=\left\{u(x_{i},y_{i})\right\}_{(x_{i},y_{i})\in\mathcal{N}^{\text{grid}}}$
@@ 2.3.1 KL expression of a Gaussian field
The $\{\xi_{m}\}_{m\in\mathbb{N}}$ in equation 6 are a set of i.i.d. standard normal variables. In this setting, the problem of estimating $\vartheta$ becomes an estimation of the realisations of the $\{\xi_{m}\}_{m\in\mathbb{N}}$ that form the desired stiffness field. For practical calculations, we will consider a finite-dimensional approximation to $g$ by truncating the series after $M$ terms. We construct the KL terms using as computational domain the unit square $D=[0,1]\times[0,1]$, which can be easily mapped to the dimensions of the beam. We model the normal field using a Matérn covariance kernel
(8) $C(x,y)=\dfrac{\sigma^{2}}{2^{\nu-1}\Gamma(\nu)}\left(\sqrt{2\nu}\dfrac{\|x-y\|_{2}}{\lambda}\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}\dfrac{\|x-y\|_{2}}{\lambda}\right)$
@@ 2.3.2 Transformation to other fields
(9) $a(x,\omega)=\mu\gamma^{-1}\left[\kappa,\dfrac{\Gamma(\kappa)}{2}\left(1+\text{erf}\left(\dfrac{g(x,\omega)}{\sqrt{2}}\right)\right)\right]$
@@
(10) $a_{\varphi}(x,\omega)=\varphi\cdot\dfrac{\exp(g(x,\omega))}{1+\exp(g(x,\omega))}+\mu\gamma^{-1}\left[\kappa,\dfrac{\Gamma(\kappa)}{2}\left(1+\text{erf}\left(\dfrac{g(x,\omega)}{\sqrt{2}}\right)\right)\right]$
@@
Figure 2: Probability density functions of random variables constructed through the transformation in Equation 10, for several values of $\varphi$. Shape and scale parameters are chosen as $\kappa=2.5$ and $\mu=0.4$, respectively.
@@ 3.1.1 Underlying idea
The multilevel Monte Carlo approach addresses the problem of prohibitively expensive forward models by exploiting a hierarchy of discretisation levels $\ell=0,\ldots L$ [19, 20]. Throughout this paper, quantities that are level-dependent will be given a subscript $\ell$. In the simple setting discussed earlier, we assumed to have a parameter $\theta$ of dimension $M$, with the PDE of Equation 4 discretized on a finite element grid with discretization parameter $R$, representing the number of finite elements on the grid. As mentioned above, we will write these parameters as $M_{\ell}$ and $R_{\ell}$ respectively, as they can change between levels. The multilevel Monte Carlo methodology allows us to estimate moments of the posterior distributions $\nu_{M_{\ell},R_{\ell}}$ for all values of $\ell$, which converge to the true underlying posterior as $M_{\ell},R_{\ell}\to\infty$ [37]. In this context, $\{R_{\ell}\}_{\ell=0}^{L}$ is an increasing sequence, where we assume, for simplicity, that there exists an $s\geq 1$ such that
@@ 3.1.3 Proposal moves
(14) $\alpha_{M_{\ell},R_{\ell}}=\min\left\{1,\dfrac{P_{M_{\ell},R_{\ell}}(\theta_{\ell}^{\prime})q_{\ell}(\theta_{\ell}^{n}|\theta_{\ell}^{\prime})}{P_{M_{\ell},R_{\ell}}(\theta_{\ell}^{n})q_{\ell}(\theta_{\ell}^{\prime}|\theta_{\ell}^{n})}\right\}$
@@
(15) $q((\theta_{\ell}^{F})^{\prime}|\theta_{\ell}^{n,F})=\theta_{\ell}^{n,F}\sqrt{1-\beta^{2}}+\beta\zeta$
@@
(16) $\alpha_{M_{\ell},R_{\ell}}=\min\left\{1,\dfrac{\mathcal{L}_{\ell}(u^{\text{obs}}|\theta_{\ell}^{\prime})\mathcal{L}_{\ell-1}(u^{\text{obs}}|\theta_{\ell}^{n,C})}{\mathcal{L}_{\ell}(u^{\text{obs}}|\theta_{\ell}^{n})\mathcal{L}_{\ell-1}(u^{\text{obs}}|(\theta_{\ell}^{C})^{\prime})}\right\}$
@@ 3.2 Incorporation of level-dependent data
The above expression is ill-suited if we would consider the limit for infinite-dimensional data [26, 35]. While a more general treatment of continuous (and thus, infinite-dimensional and spatially correlated) data is of particular interest to us, this is left for future work and the scope of the current paper will be limited to the case of high-, albeit finite-dimensional data.
@@
In order to accommodate for the increasing number of observations used between levels in the calculation of the likelihood, a weighting term should be included in the expression of the likelihood such that convergence of the vector norms of elements on increasingly fine levels is achieved. This will ensure a consistent cross-level comparison between likelihood scores, which is of vital importance as the multilevel acceptance criterion in Equation 16 contains likelihood evaluations at both levels $\ell$ and $\ell-1$.
@@
(18) $\mathcal{L}_{\ell}(u_{\ell}^{\text{obs}}|\theta_{\ell})\simeq\exp\left[\dfrac{-\|W_{\ell}(u^{\text{obs}})-F_{\ell}(\theta_{\ell})\|^{2}}{2N_{\ell}\sigma_{F}^{2}}\right]$
with $N_{\ell}$ the amount of observation points used at each level. Here $W_{\ell}:\mathbb{R}^{N_{L}}\to\mathbb{R}^{N_{\ell}}$ is a weighting function that maps all observations to the nodes used on the mesh at level $\ell$. In the case where we simply select only those observations coinciding with the coarser mesh nodes, we write $W_{\ell}=W_{\ell}^{S}$.
@@
The inclusion of $N_{\ell}$ in Equation 18 is necessary for consistency of the level-dependent likelihood expressions. Effectively, it treats the observations as if their noise level is lower at coarse levels and higher at fine levels. This can be seen as considering the observation on the coarse-level node to be a local average of all the neighbouring observations on the finest level. With an appropriate choice of $W_{\ell}$, being an actual local average, this interpretation can be made literal.
@@ 4.1 Strong convergence of PDE solution
The bound on the $\varepsilon$-cost of the multilevel MCMC algorithm depends on two aspects of the model approximation, being the truncation of the KL expansion after a finite number of terms and the finite element discretisation. The effect of the latter is a well-studied property of the multilevel Monte Carlo methodology, and the existing analyses of this carry over to our case without any alteration needed [9, 37]. To study the effect of truncating the KL expansion, the convergence analysis of [8] will be tweaked to show that similar convergence properties hold for transformations such as in Equation 10 (for the case of a Gamma random field), for any $\varphi>0$. The analysis is made in terms of the Sobolev regularity $k$ corresponding to the Sobolev space $H^{k}$. This includes Matérn covariances, for which $k=2\nu+d$ [5, 21, lord_introduction_2014].
@@ 4.1.1 Technical lemmas
The proof can be found e.g. at [17, Proposition 2.5], or [5, Theorem 3.1] or [lord_introduction_2014, Theorem 7.60] in terms of the Matérn smoothness parameter.
@@
$\|b_{m}\|_{L^{\infty}(D)}$ $\lesssim|\lambda_{m}|^{-s}$
$\|\nabla b_{m}\|_{L^{\infty}(D)}$ $\lesssim|\lambda_{m}|^{-s}$
@@
Because $C\in H^{3,0}(D^{2})$, it follows from [33, Proposition A.2] that $b_{m}\in H^{3}(D)$ for all $m\geq 1$. Let $\kappa$ be any multiindex such that $|\kappa|\in\{0,1\}$. Then it follows from the Ehrling-Nirenberg-Gagliardo inequality [25, Lemma 3.1] that
@@
The result can then be obtained analogously to the proof of [33, Theorem 2.24].
@@
The above lemma is in fact a special case of [33, Theorem 2.24], where the same bound is obtained for arbitrarily high orders of derivatives of $b_{m}$ in the case $C\in\mathcal{C}^{\infty}$. When only bounds on $b_{m}$ and its gradient are required, the smoothness requirement can be relaxed to $C\in H^{3,0}$, which enables the use of Matérn covariances with $\nu\geq 1$.
@@ 4.1.2 $L^{p}$-convergence for non-lognormal fields
The analysis of [8] shows that under assumptions $B1-B3$ a log-normal field based on a truncated KL expansion of a Gaussian field converges in $L^{p}$-sense to a log-normal field based on the true Gaussian field, and this analysis was extended by [22] to a more general setting. The next proposition shows that it can additionally be extended to Lipschitz continuous transformations of Gaussian fields that are approximated by the $\varphi$-term shown in Equation 10.
@@
$a_{\varphi}(x,\omega)$ $=\varphi\cdot\dfrac{\exp(g(x,\omega))}{1+\exp(g(x,\omega))}+a(g(x,\omega))$
@@
$\|a_{\varphi,M}-a_{\varphi}\|_{L^{p}(\mathcal{C}^{0}(\overline{D}),\Omega)}\lesssim(R_{M}^{\beta})^{1/2}$
@@
$\leq|x-y|\left(K+\varphi e^{x}+\varphi e^{y}\right)$
@@
$\|a_{\varphi,M}-a_{\varphi}\|_{L^{p}(\mathcal{C}^{0}(\overline{D}),\Omega)}\lesssim\|g_{M}-g\|_{L^{q}(\mathcal{C}^{0}(\overline{D}),\Omega)}\left\|K+\varphi e^{g_{M}}+\varphi e^{g}\right\|_{L^{r}(\mathcal{C}^{0}(\overline{D}),\Omega)}$
@@
$\|g_{M}-g\|_{L^{q}(\mathcal{C}^{0}(\overline{D}),\Omega)}\lesssim(R_{M}^{\beta})^{1/2}$
@@
$\|e^{g_{M}}\|_{L^{r}(\mathcal{C}^{0}(\overline{D}),\Omega)}\leq D_{r}$
@@
$\|a_{\varphi,M}-a_{\varphi}\|_{L^{p}(\mathcal{C}^{0}(\overline{D}),\Omega)}\lesssim(R_{M}^{\beta})^{1/2}\left(K+\varphi D_{r}+\varphi\|e^{g}\|_{L^{r}(\mathcal{C}^{0}(\overline{D}),\Omega)}\right)$

### v3

@@ 1 Introduction
(1) $u^{\text{obs}}=\mathcal{F}(\vartheta)+\eta,$
where $\eta$ represents the observational noise. The specific structure of $\eta$ impacts the observations and care should be taken when modelling this noise. The noise model used in this paper will be specified in Section 2. A common example of model (1) is when $\mathcal{F}$ is the projection onto the data space of the solution to some known partial differential equation (PDE), and $\vartheta$ are some unknown input parameters to the PDE. Within the context of this paper, we will consider a 2D plane stress problem, where the equations of linear elasticity are solved using linear finite elements (FE). These equations are introduced in Section 2.
@@ 2.1 Forward problem
(3) $\tilde{\lambda}=\dfrac{\vartheta\tilde{\nu}}{(1+\tilde{\nu})(1-2\tilde{\nu})},\qquad\tilde{\mu}=\frac{\vartheta}{2(1+\tilde{\nu})}.$
@@
(4) $-\nabla\cdot\left[\tilde{\lambda}(\nabla\cdot u)I+\tilde{\mu}(\nabla u+(\nabla u)^{T})\right]=F_{\text{body}}.$
In this equation, the term $F_{\text{body}}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ contains the body forces acting in the system, and $I$ denotes the identity matrix. Equation (4) links the spatially varying Young’s modulus to the displacement field $u:\mathbb{R}^{2}\to\mathbb{R}^{2}$. For practical calculations, we discretize this PDE on a finite element grid. We consider linear rectangular elements for the discretization of the underlying parameters [16].
@@ 2.2 High-resolution observations
(5) $\tilde{u}=\left\{u(x_{i},y_{i})\right\}_{(x_{i},y_{i})\in\mathcal{N}^{\text{grid}}},$
@@ 2.3.1 KL expression of a Gaussian field
The $\{\xi_{m}\}_{m\in\mathbb{N}}$ in equation (6) are a set of i.i.d. standard normal variables. In this setting, the problem of estimating $\vartheta$ becomes an estimation of the realisations of the $\{\xi_{m}\}_{m\in\mathbb{N}}$ that form the desired stiffness field. For practical calculations, we will consider a finite-dimensional approximation to $g$ by truncating the series after $M$ terms. We construct the KL terms using as computational domain the unit square $D=[0,1]\times[0,1]$, which can be easily mapped to the dimensions of the beam. We model the normal field using a Matérn covariance kernel
(8) $C(x,y)=\dfrac{\sigma^{2}}{2^{\nu-1}\Gamma(\nu)}\left(\sqrt{2\nu}\dfrac{\|x-y\|_{2}}{\lambda}\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}\dfrac{\|x-y\|_{2}}{\lambda}\right),$
@@ 2.3.2 Transformation to other fields
(9) $a(x,\omega)=\mu\gamma^{-1}\left[\kappa,\dfrac{\Gamma(\kappa)}{2}\left(1+\text{erf}\left(\dfrac{g(x,\omega)}{\sqrt{2}}\right)\right)\right],$
@@
(10) $a_{\varphi}(x,\omega)=\varphi\cdot\dfrac{\exp(g(x,\omega))}{1+\exp(g(x,\omega))}+\mu\gamma^{-1}\left[\kappa,\dfrac{\Gamma(\kappa)}{2}\left(1+\text{erf}\left(\dfrac{g(x,\omega)}{\sqrt{2}}\right)\right)\right],$
@@
Figure 2: Probability density functions of random variables constructed through the transformation in Equation (10), for several values of $\varphi$. Shape and scale parameters are chosen as $\kappa=2.5$ and $\mu=0.4$, respectively.
@@ 3.1.1 Underlying idea
The multilevel Monte Carlo approach addresses the problem of prohibitively expensive forward models by exploiting a hierarchy of discretisation levels $\ell=0,\ldots L$ [19, 20]. Throughout this paper, quantities that are level-dependent will be given a subscript $\ell$. In the simple setting discussed earlier, we assumed to have a parameter $\theta$ of dimension $M$, with the PDE of Equation (4) discretized on a finite element grid with discretization parameter $R$, representing the number of finite elements on the grid. As mentioned above, we will write these parameters as $M_{\ell}$ and $R_{\ell}$ respectively, as they can change between levels. The multilevel Monte Carlo methodology allows us to estimate moments of the posterior distributions $\nu_{M_{\ell},R_{\ell}}$ for all values of $\ell$, which converge to the true underlying posterior as $M_{\ell},R_{\ell}\to\infty$ [38]. In this context, $\{R_{\ell}\}_{\ell=0}^{L}$ is an increasing sequence, where we assume, for simplicity, that there exists an $s\geq 1$ such that
@@ 3.1.3 Proposal moves
(14) $\alpha_{M_{\ell},R_{\ell}}=\min\left\{1,\dfrac{P_{M_{\ell},R_{\ell}}(\theta_{\ell}^{\prime})q_{\ell}(\theta_{\ell}^{n}|\theta_{\ell}^{\prime})}{P_{M_{\ell},R_{\ell}}(\theta_{\ell}^{n})q_{\ell}(\theta_{\ell}^{\prime}|\theta_{\ell}^{n})}\right\},$
@@
(15) $q((\theta_{\ell}^{F})^{\prime}|\theta_{\ell}^{n,F})=\theta_{\ell}^{n,F}\sqrt{1-\beta^{2}}+\beta\zeta,$
@@
(16) $\alpha_{M_{\ell},R_{\ell}}=\min\left\{1,\dfrac{\mathcal{L}_{\ell}(u^{\text{obs}}|\theta_{\ell}^{\prime})\mathcal{L}_{\ell-1}(u^{\text{obs}}|\theta_{\ell}^{n,C})}{\mathcal{L}_{\ell}(u^{\text{obs}}|\theta_{\ell}^{n})\mathcal{L}_{\ell-1}(u^{\text{obs}}|(\theta_{\ell}^{C})^{\prime})}\right\}.$
@@ 3.2 Incorporation of level-dependent data
The above expression is ill-suited if we would consider the limit for infinite-dimensional data [26, 36]. While a more general treatment of continuous (and thus, infinite-dimensional and spatially correlated) data is of particular interest to us, this is left for future work and the scope of the current paper will be limited to the case of high-, albeit finite-dimensional data. To fit inside the existing multilevel MCMC framework [15], the grid resolution $R_{\ell}$ and KL truncation $M_{\ell}$ will still be allowed to enter the asymptotic regime of $\ell\to\infty$ for the convergence analysis, but $N_{L}$ is considered to reach a maximum value at finite level $L$.
@@
In order to accommodate for the increasing number of observations used between levels in the calculation of the likelihood, a weighting term should be included in the expression of the likelihood such that convergence of the vector norms of elements on increasingly fine levels is achieved. This will ensure a consistent cross-level comparison between likelihood scores, which is of vital importance as the multilevel acceptance criterion in Equation (16) contains likelihood evaluations at both levels $\ell$ and $\ell-1$.
@@
(18) $\mathcal{L}_{\ell}(u_{\ell}^{\text{obs}}|\theta_{\ell})\simeq\exp\left[\dfrac{-\|W_{\ell}(u^{\text{obs}})-F_{\ell}(\theta_{\ell})\|^{2}}{2N_{\ell}\sigma_{F}^{2}}\right],$
with $N_{\ell}$ the amount of observation points used at each level. Here $W_{\ell}:\mathbb{R}^{N_{L}}\to\mathbb{R}^{N_{\ell}}$ is a weighting function that maps all observations to the nodes used on the mesh at level $\ell$. In the case where we simply select only those observations coinciding with the coarser mesh nodes, we write $W_{\ell}=W_{\ell}^{S}$. For consistency, we additionally define $W_{\ell}$ to be the identity function on $\mathbb{R}^{N_{L}}$ for $\ell\geq L$.
@@
The inclusion of $N_{\ell}$ in Equation (18) is necessary for consistency of the level-dependent likelihood expressions. Effectively, it treats the observations as if their noise level is lower at coarse levels and higher at fine levels. This can be seen as considering the observation on the coarse-level node to be a local average of all the neighbouring observations on the finest level. With an appropriate choice of $W_{\ell}$, being an actual local average, this interpretation can be made literal.
@@ 4.1 Strong convergence of PDE solution
The bound on the $\varepsilon$-cost of the multilevel MCMC algorithm depends on two aspects of the model approximation, being the truncation of the KL expansion after a finite number of terms and the finite element discretisation. The effect of the latter is a well-studied property of the multilevel Monte Carlo methodology, and the existing analyses of this carry over to our case without any alteration needed [9, 38]. To study the effect of truncating the KL expansion, the convergence analysis of [8] will be tweaked to show that similar convergence properties hold for transformations such as in Equation (10) (for the case of a Gamma random field), for any $\varphi>0$. The analysis is made in terms of the Sobolev regularity $k$ corresponding to the Sobolev space $H^{k}$. This includes Matérn covariances, for which $k=2\nu+d$ [5, 21, 30].
@@ 4.1.1 Technical lemmas
The proof can be found e.g. at [17, Proposition 2.5], or [5, Theorem 3.1] or [30, Theorem 7.60] in terms of the Matérn smoothness parameter.
@@
$\|b_{m}\|_{L^{\infty}(D)}$ $\lesssim|\lambda_{m}|^{-s},$
$\|\nabla b_{m}\|_{L^{\infty}(D)}$ $\lesssim|\lambda_{m}|^{-s}.$
@@
Because $C\in H^{3,0}(D^{2})$, it follows from [34, Proposition A.2] that $b_{m}\in H^{3}(D)$ for all $m\geq 1$. Let $\kappa$ be any multiindex such that $|\kappa|\in\{0,1\}$. Then it follows from the Ehrling-Nirenberg-Gagliardo inequality [25, Lemma 3.1] that
@@
The result can then be obtained analogously to the proof of [34, Theorem 2.24].
@@
The above lemma is in fact a special case of [34, Theorem 2.24], where the same bound is obtained for arbitrarily high orders of derivatives of $b_{m}$ in the case $C\in\mathcal{C}^{\infty}$. When only bounds on $b_{m}$ and its gradient are required, the smoothness requirement can be relaxed to $C\in H^{3,0}$, which enables the use of Matérn covariances with $\nu\geq 1$.
@@ 4.1.2 $L^{p}$-convergence for non-lognormal fields
The analysis of [8] shows that under assumptions $B1-B3$ a log-normal field based on a truncated KL expansion of a Gaussian field converges in $L^{p}$-sense to a log-normal field based on the true Gaussian field, and this analysis was extended by [22] to a more general setting. The next proposition shows that it can additionally be extended to Lipschitz continuous transformations of Gaussian fields that are approximated by the $\varphi$-term shown in Equation (10).
@@
$a_{\varphi}(x,\omega)$ $=\varphi\cdot\dfrac{\exp(g(x,\omega))}{1+\exp(g(x,\omega))}+a(g(x,\omega)),$
@@
$\|a_{\varphi,M}-a_{\varphi}\|_{L^{p}(\mathcal{C}^{0}(\overline{D}),\Omega)}\lesssim(R_{M}^{\beta})^{1/2},$
@@
$\leq|x-y|\left(K+\varphi e^{x}+\varphi e^{y}\right).$
@@
$\|a_{\varphi,M}-a_{\varphi}\|_{L^{p}(\mathcal{C}^{0}(\overline{D}),\Omega)}\lesssim\|g_{M}-g\|_{L^{q}(\mathcal{C}^{0}(\overline{D}),\Omega)}\left\|K+\varphi e^{g_{M}}+\varphi e^{g}\right\|_{L^{r}(\mathcal{C}^{0}(\overline{D}),\Omega)}.$
@@
$\|g_{M}-g\|_{L^{q}(\mathcal{C}^{0}(\overline{D}),\Omega)}\lesssim(R_{M}^{\beta})^{1/2}.$
@@
$\|e^{g_{M}}\|_{L^{r}(\mathcal{C}^{0}(\overline{D}),\Omega)}\leq D_{r}.$
@@
$\|a_{\varphi,M}-a_{\varphi}\|_{L^{p}(\mathcal{C}^{0}(\overline{D}),\Omega)}\lesssim(R_{M}^{\beta})^{1/2}\left(K+\varphi D_{r}+\varphi\|e^{g}\|_{L^{r}(\mathcal{C}^{0}(\overline{D}),\Omega)}\right).$